OpenStudy (anonymous):

f`(x)= 5x^(2/3) - 10x^(-1/3) how this changed into 5x^(-1/3) (x-2)

OpenStudy (anonymous):

This has been factorised. Because 5 and x^(-1/3) was common to both terms it was taken out and what was left over was left in the brackets.

OpenStudy (anonymous):

how can we take common of 2/3 and -1/3 that's the point which is confusing me

OpenStudy (anonymous):

\[\frac{x^{\frac{2}{3}}}{x^{\frac{-1}{3}}}=x^{\frac{2}{3}-\frac{-1}{3}}\]\[=x^{\frac{2}{3}+\frac{1}{3}}\]\[=x^1\]\[=x\]

OpenStudy (anonymous):

You could have taken the x^(2/3) out instead but you would have ended up with x^-1.

OpenStudy (anonymous):

thanks buddy

OpenStudy (anonymous):

no worries :)